Session: Feb 2002 (session 1)
Wavelength: 18. cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (scalambra at Sun Feb 17 09:42:09 MET 2002) Success
Some interference

Torun (Kaz Borkowski at Mon Feb 18 09:54:02 MET 2002) Success
No known problems (Tsys measured normally, all BBC used worked as expected).

Jodrell_Bank (Alastair Gunn at Mon Feb 18 11:08:22 MET 2002) Minor failures
Bad count readback in VC04, therefore Tsys corrupted. Voltage still looks okay on VC04 so data should be good. No other known problems.

Cambridge (Alastair Gunn at Mon Feb 18 11:08:42 MET 2002) Failure
Cambridge did not observe due to serious fault with recorder.

Effelsberg (A. Kraus at Mon Feb 18 11:52:33 MET 2002) Success

Onsala (John Conway at Tue Feb 19 08:26:43 MET 2002) Minor failures
OK until 6 hours before end when the recorder stopped moving. Has been traced to failure of Capstan motor - EVN spare is being shipped.

Westerbork (Tony Foley at Tue Feb 19 14:37:21 MET 2002) Minor failures
Scans 07:43 to 08:05 out of HA range